OpenStudy (anonymous):

I am needing help with finding an integrating factor and solving a given equation. For problems such as: y + (2xy - e^(-2y))y' = 0

OpenStudy (anonymous):

Not a linear equation, so I assume you mean the integrating factor an (in)exact equation. \[y+(2xy-e^{-2y})y'=0\] Here, \(M(x,y)=y\) and \(N(x,y)=2xy-e^{-2y}\). \[\frac{\partial M}{\partial y}=1~~~~~~~~~\frac{\partial N}{\partial x}=2y\] Clearly not equal. The integrating factor is given by \[\mu(y)={\large e}^{\displaystyle\int\frac{2y-1}{y}~dy}=e^{2y-\ln y}=\frac{e^{2y}}{y}\] \[\frac{e^{2y}}{y}\cdot y+\frac{e^{2y}}{y}(2xy-e^{-2y})y'=0\\ e^{2y}+\left(2xe^{2y}-\frac{1}{y}\right)y'=0\] Equation is now exact, but just to check, let \(M=e^{2y}\) and \(N=2xe^{2y}-\dfrac{1}{y}\). Then, \[\frac{\partial M}{\partial y}=2e^{2y}~~~~~~~~~\frac{\partial N}{\partial x}=2e^{2y}\]

OpenStudy (anonymous):

Thank you! That was exactly what I needed. Your answer will help me apply this to the remaining questions in the section!